what is a magnitude of the electric field at a point midway between a -6,5uC and +5.4uC charge 8.0cm apart?
1
Expert's answer
2021-10-19T09:45:14-0400

"E_1=\frac{kq_1}{r^2}\\E_1=\frac{{k}\times{(-6.5\times10^{-6})}}{4^2\times10^{-4}}"

"E_1=-3.656\times10^7N/c"

"E_2=\frac{kq_2}{r^2}\\E_2=\frac{{k}\times{(5.4\times10^{-6})}}{4^2\times10^{-4}}"

"E_2=3.037\times10^7N/c"

"E=|E_1|+|E_2|"

"E=6.693\times10^{7}N/c"